In a population, the melanic, black-colored seagull, is due to a dominant allele. Mm and MM individuals are black, and mm individuals are white. In each population, 81 individuals are white-colored, with 19 black- colored individuals. Assuming HWE, what is the approximate frequency of the melanic (M) allele in the population?

In class, we took the square root of .81 to get the frequency of the recessive allele, but I don't know how that could be the frequency of the recessive allele since we did not take into consideration the little m that is present in the heterozygote. Can you please explain how we were able to ignore the recessive freq in the heterozygote
Answer #1

White seagull = mm = 81

Genotypic frequency of white seagulls = 81/100 = 0.81 = mm

Allelic frequency of white seagulls = m = √0.81 = 0.9

Allelic frequency of black seagulls = M = 1 - 0.9 = 0.1

As per HWE,

p (dominant allele frequency) + q (recessive allele frequency) = 1

Genotypic frequency of heterozygotes = 2pq = 2Mn = 2×0.1×0.9 = 0.18

Genotypic frequency of homozygote dominant = pp = MM = 0.1×0.1 = 0.01

2pq + pp = 0.19 = frequency of black seagull

All should be = 1

Here, frequency of black seagull (0.19) + frequency of white seagull (0.81) = 1

Please rate.